OpenStudy (anonymous):

simplify: (3x/y) * (3x^2y^-2/12y^3)

OpenStudy (anonymous):

@ChmE

OpenStudy (anonymous):

\[\frac{ 3x }{ y }*\frac{ 3x^2 }{ 12y^5 }=\frac{ 3x^3 }{4y^6 }\]

OpenStudy (anonymous):

Can you show me how you simplified that

OpenStudy (anonymous):

\[({3x \over y})({3x^2y^{-2} \over 12y^3})\]\[x^ax^b=x^{a+b}\]\[{x^a \over x^b}=x^{a-b}\]Memorize these

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

We'll do one part at a time \[3x \times 3x^2=9x^3\]any questions on this part

OpenStudy (anonymous):

no you can continue!

OpenStudy (anonymous):

I want you to do the multiplication of the denominator

OpenStudy (anonymous):

do you find you LCD?

OpenStudy (anonymous):

Multiplication of fractions you just simply multiply across

OpenStudy (anonymous):

\[\frac{ a }{ b } \times \frac{ c }{ d }=\frac{ ac }{ bd }\]

OpenStudy (anonymous):

oh okay I thought you needed to find your common denominator first.

OpenStudy (anonymous):

So what is this\[y \times 12y^3\]

OpenStudy (anonymous):

12y^4

OpenStudy (anonymous):

yup now we have to simplify the y's in the numerator and the denominator.\[\frac{ y^{-2} }{ y^4 }=y^{-2-4}=y^{-6}=\frac{ 1 }{ y^6 }\]

OpenStudy (anonymous):

Okay I'm understanding but we also have our 9/12= 3/4

OpenStudy (anonymous):

Now we are left with \[\frac{ 3x }{ y }*\frac{ 3x^2 }{ 12y^5 }=\frac{ 9x^3 }{12y^6 }\]Just simplify the constants and we get the answer the other guy posted

OpenStudy (anonymous):

oh yeah I forgot about our x^3 \[3x ^{3}/4y ^{6}\]